OpenStudy (anonymous):

Determine the point on the curve y=x^2 which is nearest point (8, 0)

OpenStudy (anonymous):

well we know the shortest distance between a point and a line is a straight line whose slope is the opposite reciprocal of the given function's slope....Which we can find by taking the derivative of the function. y'=2x So the slope of the line is we seek is -1/2... and our given point is (8,0) Hope that just about gives you what you need to know!

OpenStudy (anonymous):

I need to find a point, not a line though

OpenStudy (amistre64):

the nearest point is perp to the curve

OpenStudy (amistre64):

you need a point with a slope of -1/2x that hits the point (8,0)

OpenStudy (amistre64):

or better yet, prolly define a generic point that slopes out correctly with (8,0)

OpenStudy (amistre64):

given points (x,x^2) and (8,0) slope is defined as: x^2 ----- x-8 and this needs to equal -1/2x

OpenStudy (amistre64):

put it all together and solve for x :)

OpenStudy (anonymous):

okay that makes sense.. thanks

OpenStudy (amistre64):

you might get an extra value so chk both results

OpenStudy (anonymous):

I got 2x^3+x-8=0

OpenStudy (anonymous):

how do I solve that?

OpenStudy (amistre64):

carefully :)

OpenStudy (amistre64):

thats gonna have min max at 6x^2 + 1 = 0; so at x = +- sqrt(1/6) which can narrow down the domain

OpenStudy (anonymous):

how am I supposed to find the y coordinate from there?

OpenStudy (amistre64):

a cubic looks like this:so knowing where the bends are can narrow down the interval we are looking for is my idea for that

OpenStudy (amistre64):

when x=0 we have a change of concavity, so its a good assumption that x is going to rott out between sqrt(1/6) and 8

OpenStudy (amistre64):

id start with 1 and try to work out a few integers to see if you can hit a good integer

OpenStudy (amistre64):

2+1-8 is negative 2(2)^3+2-8=0 16 - 6 is postitive, so it crosses someplace between 1 and 2

OpenStudy (anonymous):

I think this is a lot more difficult than it needs to be

OpenStudy (amistre64):

prolly, but i have to rely upon my wits since I aint got your material to guide me :)

OpenStudy (anonymous):

if you start from the beginning, the point I'm looking for is (x, x^2) so couldn't I use the distance formula and set it equal to zero then find the derivative to solve for x?

OpenStudy (amistre64):

you dont want a distance of zero; 8,0 aint on the curve

OpenStudy (anonymous):

but when you're finding max and mins with derivatives, don't you want to set the derivative equal to zero?

OpenStudy (amistre64):

hmmm, so your idea is to find the min distance with a derivative of the distance formula? might work ...

OpenStudy (anonymous):

Yeah, I vaguely remember someone mentioning that in class last week

OpenStudy (amistre64):

id have to remember back many years to recall someone mentioning it in class lol

OpenStudy (anonymous):

to find the derivative of a sqrt you would have to use the chain rule, right?

OpenStudy (amistre64):

if you derive the distance formula you only gotta worry about the derivative of the innards that make up the numerator

OpenStudy (amistre64):

chain rule would work; yes

OpenStudy (amistre64):

sqrt(x) derives to x'/2sqrt(x)

OpenStudy (amistre64):

and a fraction is only zero when the numerator is zero; so focus on the derivative of the innards

OpenStudy (amistre64):

x^2^2 +(x-8)^2 4x^3 +2(x-8) = 2(2x^3+x-8) lol same thing

OpenStudy (anonymous):

yeah so would my answer just be (-sqrt(1/6), 1/6)

OpenStudy (amistre64):

no, that is just a bending point; the solution is in the interval between 1 and 2

OpenStudy (amistre64):

we can do a newton rhapsody method to narrow it closer and closer if need be

OpenStudy (anonymous):

I don't understand...